# Zonal vs. Nodal Pricing: An Analysis of Different Pricing Rules in the German Day-Ahead Market

### v4

@@ Abstract
The European electricity market is based on large pricing zones with a uniform day-ahead price. The energy transition leads to changes in supply and demand and increasing redispatch costs. In an attempt to ensure efficient market clearing and congestion management, the EU Commission has mandated the Bidding Zone Review (BZR) to reevaluate the configuration of European bidding zones. Based on a unique data set published in the context of the BZR for the target year 2025, we compare various pricing rules for the German power market. We compare market clearing and pricing for different zonal and nodal models, including their generation costs and associated redispatch costs. In numerical experiments with this dataset, the differences in the average prices in different zones are low. Congestion arises as well, but not necessarily on the cross-zonal interconnectors. The total costs across different configurations are similar and the reduction of standard deviations in prices is also small. This might be different with other load and generation scenarios, but the BZR data is important as it was created to make a decision about splits of the existing bidding zones. Nodal pricing rules lead to the lowest total cost. We also evaluate differences of nodal pricing rules with respect to the necessary uplift payments, which is relevant in the context of the current discussion on non-uniform pricing in the EU. While the study focuses on Germany, the analysis is relevant beyond and feeds into the broader discussion about pricing rules in non-convex markets.
@@
journal: Operations Research
@@ 1 Introduction
A central discussion on electricity market pricing revolves around zonal and nodal pricing (aka. locational marginal pricing). Among the electricity wholesale markets implementing nodal pricing are Argentina, Chile, Mexico, New Zealand, Peru, Russia, Singapore, and several regions in the United States (U.S.). In contrast, a zonal pricing approach has been employed in the large and coupled EU market. Large groups of nodes are aggregated into zones, and the market clearing problem only considers flow constraints between zones. It is supposed that no or only limited congestion occurs inside a zone. With some exceptions (e.g., Italy, Norway, Sweden), there is a single price zone within a country, and the day-ahead market provides a uniform national electricity price. Ignoring transmission constraints can lead to dispatch decisions that are not physically feasible. Thus, after the market clearing, transmission operators conduct redispatch to ensure that the final allocation aligns with physically feasible power flows. In the USA, all Independent Service Operators (ISOs) moved from a zonal to a nodal model, partly due to steeply increasing redispatch costs. Aravena et al., (2021) provide an excellent summary of historical developments in different juristictions.
@@
Due to climate change and the ongoing energy transition, the zonal pricing paradigm has further come under scrutiny (Eicke & Schittekatte,, 2022; Bertsch et al.,, 2016; Trepper et al.,, 2015). Volatile renewable energy sources such as wind and solar lead to congestion in different parts of the electricity grid (Neuhoff et al.,, 2013). In addition, the energy transition can lead to structural problems. We will focus on Germany as a case in point. Energy is cleared as if there were no transmission constraints in Germany, but the traded energy can often not be delivered due to transmission constraints. As a result, redispatch volumes and costs have increased substantially in recent years to 34 TWh and EUR 3.1 billion in 2023 (Bundesnetzagentur,, 2024). In an attempt to ensure efficient market clearing and congestion management, the EU Commission – as part of the Clean Energy Package – has mandated a Bidding Zone Review (BZR) to reevaluate the configuration of European bidding zones. An integral part of this process is a locational marginal pricing (LMP) study conducted by the European Network of Transmission System Operators for Electricity (ENTSO-E) as a basis to identify structural congestion and modified bidding zones. Having accurate data and grid models is central for the quality of the BZR. The LMP study constitutes the results of a multi-year effort of the European TSOs to provide a representative data set, and it is unparalleled in terms of the quality and degree of detail of the data. The study is based on an up-to-date model of the electricity grid and supply and demand estimates for 2025 and different climate scenarios. In this study, ENTSO-E solved linearized unit commitment models with marginal pricing to obtain nodal prices (ENTSO-E,, 2022). Based on the resulting prices, in August 2022, the European Union Agency for the Cooperation of Energy Regulators (ACER) decided on a set of alternative bidding zone configurations (ACER, 2022a, ). Following Article 16 of the BZR methodology, ENTSO-E has published non-confidential data related to the LMP study (ENTSO-E,, 2023). We refer to this data set as the BZR data in what follows and leverage this dataset to expand on the results of the LMP study, focusing on the German bidding zone. Germany is particularly interesting, because ACER suggested the evaluation of several configurations with splits of the large German-Luxembourgish bidding zone.
@@ 1.2 Related Literature
A few recent studies aimed to analyze the impact of a bidding zone split on Germany or other countries in Europe. These studies are typically based on proprietary data sets and models and they assume different target years. The consulting company Aurora published results of another study where they predict 5 Euros per MWh price difference between the north and the south of Germany in the years 2030, increasing to 9 Euros in 2045.111https://auroraer.com/insight/power-market-impact-of-splitting-the-german-bidding-zone/ Models and data are proprietary. In another study by the EWI and THEMA consulting [^2] https://www.ewi.uni-koeln.de/en/publications/price-impact-of-a-german-bidding-zone-split/ the authors predict differences in prices in a southern and a northern zone of around 15 to 10 Euros per MWh between 2024 and 2033, with a decreasing trend. The study is based on two separate models and data by the EWI and THEMA consulting that are not publicly available. The size of the spread differs in THEMA’s and EWI’s results according to the authors. For 2030 the study makes a number of assumptions regarding electrolysers and new lines being built, batteries, and a shift of renewables from North to South. Tiedemann et al., (2024) discuss the market value of renewable energy sources after a zonal split for 2030. Their study is based on a data set and grid model by the Fraunhofer IEE and they predict around 10 Euros per MWh difference between the north and south of Germany respectively in 2030. They do not comment on reductions in redispatch by a zonal split.
@@
It is worth mentioning that the data sets used in these studies and the underlying grid models differ substantially, and so do the target years analyzed. For example, there are differences in the number of lines in the BZR data set compared to the widely used JAO Static Grid model. [^3] https://www.jao.eu/static-grid-model,accessedinMay2024. Also, the assumptions made for supply and demand for the future differ. We stick to the BZR dataset which appears to have a very comprehensive grid model and a near-term target year of 2025.444There are also earlier studies on a zonal split, such as those by Trepper et al., (2015), but the data sets and assumptions are even harder to compare due to the significant changes in the past ten years.
@@
All of these studies analyze static models where demand and supply are given as we do. Ambrosius et al., (2020) use a multi-level mixed-integer non-linear model based on Grimm et al., (2019) to analyze investment incentives over time. Such models cannot accommodate the level of detail in static models due to their computational complexity, such that the paper is based on an aggregation of the German electricity network into 28 nodes. However, the approach complements static models taking into account the long-run effects of prices on investment. Several articles discuss the impact of price splits on transmission and generation investment incentives and emission reductions (Grimm et al.,, 2016, 2021, 2022). For example, Grimm et al., (2021) predict that a split of the German market into two price zones has only a small impact on efficiency.
Another stream in the literature asks how zones should be split. While some papers focus on identifying configurations by clustering power transfer distribution factors (PTDFs) (e.g., (Kumar et al.,, 2004) and (Kłos et al.,, 2014)), most works use LMPs as clustering features (Stoft,, 1997). Burstedde, (2012) computes LMPs for the scenario years 2015 and 2020 considering a simplified 72-node representation of the European grid, and evaluates the shapes and sizes of the zones obtained by clustering these LMPs using hierarchical clustering. Breuer et al., (2013) implement a genetic algorithm and apply it on a dataset consisting of 380 kV nodes considering a projected system for 2016 and 2018. Based on six scenarios, Felling & Weber, (2018) compute robust price zone configurations for Central Western Europe by employing hierarchical clustering based on the LMPs of approximately 2,200 nodes and assigning weights to nodes depending on their demand- and supply situation. In a complementary paper, Dobos et al., (2024) use the BZR data and the published nodal prices to compute clusters with low price variance. They show that the clusterings they found are not robust across time or the clustering method used.
@@
Finally, there is an extensive literature on nodal pricing and different pricing rules to deal with the non-convexities of power markets. We refer the interested reader to Liberopoulos & Andrianesis, (2016) for an excellent overview. The most prominent examples are Integer Programming (IP) pricing (O’Neill et al.,, 2005) and Convex Hull (CH) pricing (Gribik et al.,, 2007; Hogan & Ring,, 2003; Bichler et al.,, 2022). IP pricing is used in a number of U.S. markets and so are versions of CH pricing. Both pricing rules require side-payments by the market operator to compensate losses of market participants (make-whole payments, MWPs) or even to compensate all incentives to deviate from the efficient allocation (lost opportunity costs, LOCs). These pricing rules are now being discussed in Europe in the context of non-uniform pricing rules. CH pricing is known to provide poor congestion signals. We also draw on a recent proposal by Ahunbay et al., (2024) that trades off make-whole payments and the quality of the congestion signal and evaluate these three pricing rules in the context of our study, which provides input for the discussion on non-uniform pricing in the EU day-ahead market (All NEMO Committee,, 2023).
@@ 1.3 Contributions
Our goal is to understand how prices in these configurations compare to nodal prices and the current market design with a uniform national price. Specifically, we compare the total costs and prices that arise for national, zonal, and different nodal models. For this we leverage the BZR data set, which provides the most comprehensive and most recent data set available about the European power grid and serves as a foundation for decisions by EU authorities.
In contrast to earlier studies, the average price differences across zonal configurations between 1 and 4 zones for Germany are low, less than 3.5 EUR/MWh. There are also individual days where the price difference between zones is more than 10 EUR/MWh, but on average the differences are small. For the scenarios in the BZR data with climate year 2009, the excess supply in the north of Germany after computing a nodal dispatch was typically such that the cross-zonal lines or interconnectors were not a binding contraint. There was congestion in the network, but not necessarily on the interconnectors. [^5] In 2018, the Austria was split from the joint price zone with Germany. Demand and supply is significantly different between both countries, yet prices (after the gas crises in 2021 and 22) are often remarkably close as well (see https://www.energy-charts.info/). This finding extends to differences in the redispatch between different zonal configurations. It is difficult to model real-world redispatch practices of TSOs. We could only implement two simplified minimal-cost and minimal-volume redispatch methods. Real-world practices are more complicated, but we argue that it provides a reasonable lower bound for redispatch costs. While these redispatch costs decrease slightly from one to two zones, we see little effect with further splits. The cross-zonal capacities play a role for these differences and we will discuss them in more detail below. Our findings do not mean that there cannot be scenarios of supply and demand in Germany where cross-zonal links become binding and there are differences in prices. However, in the BZR dataset, intended to inform ACER and other stakeholders, this is not what we find.
The BZR dataset also provides us with an opportunity for an apples-to-apples comparison of zonal and nodal pricing rules. So, we can quantify the effect of moving from a zonal to a nodal pricing regime. In our study, nodal prices lead to higher generation costs than zonal pricing, but to lower total costs taking our estimated redispatch costs into account. This is due to the fact that redispatch is largely avoided in nodal pricing systems. However, there are different ways how nodal prices can be computed and these nodal pricing rules are currently being discussed in the context of the new capacity management rules CACM 2.0 (All NEMO Committee,, 2023). In other words, the comparison between nodal and zonal prices also depends on the nodal pricing rule. We consider Integer Programming (IP) pricing (O’Neill et al.,, 2005) (as used in U.S. power markets) and Convex Hull (CH) pricing (Hogan & Ring,, 2003; Gribik et al.,, 2007) as two established pricing rules in U.S. markets, and the recently proposed Join pricing rule (Ahunbay et al.,, 2024). The results illustrate the trade-offs that can be expected in terms of quality of the congestion signals and minimum make-whole payments. [^6] We compared the nodal prices that we computed to those published in the BZR dataset. The latter were based on duals of a linear programming relaxation of entire European dispatch problem. Details of this model have not been released but the average nodal prices that we computed are not significantly different to those pubished in the BZR dataset in a sample of hourly prices analyzed. While our analysis focuses on Germany, the policy implications of our study go beyond in that they analyze the differences between different zonal and nodal pricing rules based on a realistic dataset.

### v6

@@ Abstract
The European electricity market is based on large pricing zones with a uniform day-ahead price. The energy transition leads to changes in supply and demand and increasing redispatch costs. In an attempt to ensure efficient market clearing and congestion management, the EU Commission has mandated the Bidding Zone Review (BZR) to reevaluate the configuration of European bidding zones. Based on a unique data set published in the context of the BZR for the target year 2025, we analyze the short-run effects of various pricing rules for the German-Luxembourgish bidding zone. We compare market clearing and pricing for different zonal models, including their generation and redispatch costs. In numerical experiments with this dataset, the differences in the average prices in different zones are low. The total costs across different configurations are similar and the reduction of standard deviations in prices is also small. This might be different with other load and generation scenarios, but the BZR data is important as it was created to make a decision about splits of the existing bidding zones. We can replicate several results from the BZR study, except the large cost savings when moving from one to two price zones in Germany and Luxembourg. In addition to the four zonal configurations analyzed in the BZR study, we compare these against a nodal pricing system. While the total cost savings after introducing zonal splits were less than 1%, nodal pricing led to savings of 5-6%. We also evaluate differences of nodal pricing rules with respect to the necessary uplift payments, which is relevant in the context of the discussion on non-uniform pricing in the EU. While the study focuses on Germany, the analysis is relevant beyond and feeds into the broader discussion about pricing rules in non-convex markets.
@@
journal: Energy Economics
@@ 1 Introduction
A central discussion on electricity market pricing revolves around zonal and nodal pricing (aka. locational marginal pricing). Among the electricity wholesale markets implementing nodal pricing are Argentina, Mexico, New Zealand, Russia, Singapore, and several regions in the United States (U.S.). In contrast, a zonal pricing approach has been employed in the large and coupled EU market. Large groups of nodes are aggregated into zones, and the market clearing problem only considers flow constraints between zones. It is supposed that no or only limited congestion occurs inside a zone. With some exceptions (e.g., Italy, Norway), there is a single price zone within a country, and the day-ahead market provides a uniform national electricity price. Ignoring transmission constraints can lead to dispatch decisions that are not physically feasible. Thus, after the market clearing, transmission operators conduct redispatch to ensure that the final dispatch solution (or, allocation) satisfies the power flow equations. In the USA, all Independent System Operators (ISOs) moved from a zonal to a nodal model, partly due to steeply increasing redispatch costs. Aravena et al., (2021) provide an excellent summary of historical developments in different jurisdictions.
@@
Due to the ongoing energy transition, the zonal pricing paradigm has further come under scrutiny (Eicke & Schittekatte, , 2022; Bertsch et al., , 2016; Trepper et al., 2015a, ). Volatile renewable energy sources such as wind and solar, combined with the limitations of the electricity grid can lead to system overloads – commonly referred to as congestions or bottlenecks (Neuhoff et al., , 2013). Structural congestion refers to predictable and recurring bottlenecks. We will focus on Germany as a case in point. In this region, energy is cleared as if there were no transmission constraints, but the traded energy can often not be delivered due to transmission constraints. As a result, redispatch volumes and costs occur, which have increased substantially in recent years to 34 TWh and EUR 3.1 billion in 2023 (Bundesnetzagentur, , 2024). In an attempt to ensure efficient congestion management, the EU Commission – as part of the Clean Energy Package – has mandated a Bidding Zone Review (BZR) to reevaluate the configuration of European bidding zones. An integral part of this process is a locational marginal pricing (LMP) study conducted by the European Network of Transmission System Operators for Electricity (ENTSO-E) as a basis to identify structural congestion and modified bidding zones. Having accurate data and grid models is central for the quality of the BZR. The LMP study constitutes the results of a multi-year effort of the European TSOs to provide a representative data set, and it is unparalleled in terms of the quality and degree of detail of the data. This data set includes an electricity grid model derived from the Ten-Year Network Development Plan (TYNDP) 2020111Ongoing network expansion projects such as the development of the Südlink connection (expected to be finished in 2028) are not modeled in the data set., and supply and demand projections for 2025 under three climate years. In the LMP study, ENTSO-E solved linearized unit commitment models with marginal pricing to obtain nodal prices (ENTSO-E, , 2022). Based on these prices, in August 2022, the European Union Agency for the Cooperation of Energy Regulators (ACER) decided on a set of alternative bidding zone configurations (ACER, 2022b, ). Following Article 16 of the BZR methodology, ENTSO-E has published non-confidential data related to the LMP study (ENTSO-E, , 2023). In what follows, we refer to this data set as the BZR data and leverage it to expand on the results of the LMP study, focusing on the German BZ. [^2] Although Germany and Luxembourg together form a single BZ, we will refer to it simply as the German BZ. This bidding zone is particularly interesting, because ACER suggested four alternative configurations (ACER, 2022c, ) that define a split of the large zone into 2, 3 or 4 smaller zones. [^3] The configurations were identified by solving a clustering problem, where the features represent the time series of prices computed in the LMP study (ACER, 2022a, ).. Subsequently, the German TSOs proposed a new configuration comprising 5 zones, with the Schleswig-Holstein region represented as a separate zone.
1.2 Challenges in Congestion Management
A key driver of the BZR is ensuring efficient congestion management. Congestion occurs when transmission capacity is insufficient to accommodate the flow of electricity from generation units to demand centers, leading to bottlenecks that require operational adjustments like redispatch or curtailment. Over the last years, redispatch costs and curtailment rates have increased considerably in Europe due to the increasing integration of renewable energy sources (RES), limited storage capabilities and insufficient transmission capacity. In 2024, the European Commission proposed an emissions reduction of 90% by 2040 compared to the 1990 levels (European Commission, , 2024). This objective depends on the deployment of RES. Current efforts primarily focus on installing renewable capacity in areas with the highest resource potential and overlook the grid topology and the local demand. Regions rich in renewable generation potential may not align with areas of high electricity demand. Thus, the need to transmit power within zones could frequently surpass the capacity of the grid, leading to congestion and increased curtailment of RES. Newbery, (2023) and Newbery & Biggar, (2024) discuss the inefficiency of current European market designs from the perspective of the RES merchants, emphasizing that investments are driven by average curtailment rates, which fail to account for significantly higher marginal curtailment rates. [^4] Marginal curtailment represents the extra curtailment caused by the last MW of capacity installed. Average curtailment is defined as the total curtailed volume divided by the total installed capacity. This misalignment risks promoting overinvestment in poorly connected areas, since, according to the authors, marginal curtailment is more than three times higher than average curtailment. A recent study by Thomassen et al., (2024) suggests that in a grid expansion scenario that extrapolates current trends, up to 310 TWh of renewable generation could be curtailed in 2040 due to the limitations in the grid. Moreover, even in an extreme grid expansion scenario that involves expanding the length of the European grid by more than a third, the total redispatch volumes would increase almost sixfold compared to 2022. While the increase in curtailment rates and redispatch volumes is predictable, addressing these challenges requires a thorough evaluation of alternative market designs and pricing techniques.
@@ 1.3 Related Literature
A few recent studies aimed to analyze the impact of a bidding zone split on Germany or other countries in Europe. These studies are typically based on proprietary data sets and models, and assume different target years. The consulting company Aurora published results of such a study where they predict 5 EUR/MWh price difference between the north and the south of Germany in the year 2030, increasing to 9 EUR in 2045 (Aurora, , 2023). Models and data are proprietary. Another study by EWI and THEMA consulting (Diers et al., , 2023) predicts differences in prices in a southern and a northern zone of around 10-15 EUR/MWh between 2024 and 2033, with a decreasing trend. The study is based on two separate models and data by EWI and THEMA consulting that are not publicly available. The size of the spread differs in THEMA’s and EWI’s results according to the authors. For 2030, the study makes a number of assumptions regarding electrolysers and new lines being built, batteries, and a shift of renewables from north to south. Tiedemann et al., (2024) discuss the market value of renewable energy sources after a zonal split for 2030. Their study is based on a data set and grid model by Fraunhofer IEE and they predict around 10 EUR/MWh difference between the north and south of Germany in 2030. They do not comment on reductions in redispatch. Moreover, Gerhardt et al., (2025) show that a a locally organized electricity market increases the cost efficiency of a climate-neutral electricity system.
@@
It is worth mentioning that the data sets used in these studies and the underlying grid models differ, and so do the target years analyzed. For example, there are differences in the number of lines in the BZR data set compared to the widely used JAO Static Grid model. [^5] JAO Static Grid model contains 837 lines $\geq$ 220kV (see https://www.jao.eu/static-grid-model, accessed in May 2024). The dataset we use consists of 2407 lines spanning various voltage levels, both below and above 220kV. Also, the assumptions made for the future supply and demand and the redispatch methods used in these studies differ. [^6] There are also earlier studies on a zonal split, such as those by Trepper et al., 2015a , but the data sets and assumptions are even harder to compare due to the significant changes in the past ten years. The study by EWI and THEMA, while not disclosing the implementation details, calculated redispatch volumes “using a welfare optimal redispatch procedure with limited countertrading”. The study by Aurora does not disclose any details regarding how redispatch is computed.
In November 2024, the average market prices for the alternative configurations were published (ENTSO-E, , 2024). Importantly, the BZR study evaluates the alternative configurations at the Central European (CE) level, whereas our study evaluates them at the German level. We highlight the key modelling differences between our study and the BZR study in E. The BZR results reveal that for the scenario year 2009 (see Section 2.2), a two-zone split leads to a zonal price difference of 4.77 EUR/MWh. Averaging over all climate years considered in the BZR, the average prices observed in the north and in the south differ by $\approx$ 7 EUR/MWh. The results published by ENTSO-E also suggest that a split of Germany into 2-5 zones leads to price increases higher than 1 EUR/MWh in Czechia, Austria, Hungary, Slovakia, Slovenia, Croatia and Romania. The final results of the BZR study were published in April 2025 (ENTSO-E, , 2025) and show that the welfare gains associated with any alternative configuration are less than 1% of the simulated system costs in the CE region. These configurations also lead to redispatch savings of $\approx$ 50% compared to the status quo European BZ layout. We analyze these results in detail in F.
@@
All of the previously mentioned studies analyze static models where demand and supply are given as we do. Ambrosius et al., (2020) use a multi-level mixed-integer non-linear model based on Grimm et al., (2019) to analyze investment incentives over time. Such models cannot accommodate the level of detail in static models due to their computational complexity, such that the paper is based on an aggregation of the German electricity network into 28 nodes. However, the approach complements static models taking into account the long-run effects of prices on investment. Several articles discuss the impact of price splits on transmission and generation investment incentives and emission reductions (Grimm et al., , 2016, 2021, 2022). For example, Grimm et al., (2021) predict that a split of the German market into two price zones has only a small impact on efficiency. Considering 3 grid expansion scenarios to be implemented by 2040, Thomassen et al., (2024) analyze the evolution of redispatch costs for a large part of the EU based on a model consisting of 1024 nodes. Their study is built upon the PyPSA-Eur framework (Hörsch et al., , 2018) and shows that the redispatch volumes and costs will increase substantially regardless of the grid expansion scenario. Several studies, focusing on various jurisdictions, estimate the static dispatch inefficiencies of a zonal market design to range between 0.5% and 3% (Simshauser, , 2025; Holmberg & Tangerås, , 2023; Green, , 2007; Leuthold et al., , 2008; Aravena & Papavasiliou, , 2016).
Another stream in the literature asks how zones should be split. Most studies use heuristics to identify price zones configurations by clustering power transfer distribution factors (Kumar et al., , 2004; Kłos et al., , 2014) or LMPs (Stoft, , 1997; Burstedde, , 2012; Breuer et al., , 2013; Felling & Weber, , 2018). In contrast, Grimm et al., (2019) propose a mixed-integer nonlinear model to divide a market area into a specific number of price zones, ensuring the resulting configuration is welfare-optimal. Ambrosius et al., (2020) extend this model by including capacity investment decisions, aiming to determine the available transfer capacities between zones. Dobos et al., (2025) use the BZR data and the published nodal prices to compute zones with low price variance. They show that the clusterings they found are not robust across time or clustering method used.
@@
Finally, there is an extensive literature on nodal pricing and different pricing rules to deal with the non-convexities of power markets Liberopoulos & Andrianesis, (2016). The most prominent examples of nodal pricing rules are Integer Programming (IP) pricing (O’Neill et al., , 2005) and Convex Hull (CH) pricing (Gribik et al., , 2007; Hogan & Ring, , 2003; Bichler et al., , 2022). Both are used in a number of U.S. markets (MISO, , 2023; PJM, , 2018) and require side-payments by the market operator to compensate losses of market participants (make-whole payments) or even to compensate all incentives to deviate from the efficient allocation (lost opportunity costs). These pricing rules are now being discussed in Europe in the context of non-uniform pricing rules. In particular, Pollitt, (2023) analyzes whether and how nodal prices should be implemented in Europe. CH pricing is known to provide poor congestion signals. We also draw on a recent proposal by Ahunbay et al., (2024) that trades off make-whole payments and the quality of congestion signals, and evaluate these three pricing rules in the context of our study, which provides input for the discussion on non-uniform pricing in the EU day-ahead market (All NEMO Committee, , 2023).
@@ 1.4 Contributions
In this study, we conduct an independent analysis of the BZR dataset, which is based on supply and load data collected in 2019 and models the target year 2025. This dataset is arguably the most comprehensive and recent dataset available for the European power grid and serves as a foundation for decisions by the European Commission. Unlike the BZR study, which simulates the entire CE continent, we focus exclusively on modeling Germany and its surrounding nodes, which allows for exact optimization of the dispatch and price computation and a focus on the effects in Germany.
Our analysis considers not only zonal but also nodal pricing, for which we evaluate the average costs and day-ahead prices. For the nodal market design, we formulate the unit commitment and economic dispatch problem, or welfare maximization problem, as a mixed-integer linear program (MILP). This differs from the LMP study, which used a linear programming relaxation of the problem for CE due to its large size (ENTSO-E, , 2022). We then compute prices using three distinct pricing approaches: IP (Integer Programming) pricing (as used in U.S. power markets) (O’Neill et al., , 2005), CH (Convex Hull) pricing (Hogan & Ring, , 2003; Gribik et al., , 2007), and Join pricing (Ahunbay et al., , 2022). These methods compute prices based on the optimal allocation solution, minimizing different classes of lost opportunity costs. We note the difference to the LMP study, in which nodal prices were computed as shadow prices for the relaxed allocation problem. While the nodal prices computed in our study are close to the prices computed in the LMP study on average, they can differ on individual nodes. For the climate year 2009, the average value of the prices in the LMP study is 46.83 EUR/MWh, while the average value of the prices computed in this study with IP pricing is 46.63 EUR/MWh.
For the zonal market design, we evaluate Germany’s current single-zone configuration and the alternative configurations proposed by ACER, which include 2, 3, or 4 zones (ACER, 2022c, ). We formulate the zonal allocation problem as an MILP and model the cross-zonal capacities using a net transfer capacity approach with security limits. In a subsequent step, we compute zonal prices using the three methods mentioned previously and analyze the redispatch costs associated with each configuration. These costs are identified based on a minimum-cost redispatch procedure, which assumes that all generators are redispatchable and that redispatch is determined through a system-wide optimization, i.e., there is a cross-border cooperation between German TSOs, but not across all states in CE. Importantly, while the implementation details of the redispatch algorithm used in the BZR study are not published, the algorithm assumes full coordination in CE. Such redispatch process would mean that the most efficient remedial action within the whole CE region is selected to relieve a bottleneck in the German grid, which is against current practice.
1.5 Key Results
Let us now summarize our main findings with respect to prices, price differences between zones, total cost savings due to the zonal configurations and savings in redispatch costs.
As indicated earlier, the LMP prices that we computed are very close to the prices published in the BZR study (the difference between the average prices in both studies is less than 1 EUR/MWh), which confirms their findings. In contrast to earlier studies, the average price differences across zonal configurations between 1 and 4 zones for Germany are low in our study, less than 3.5 EUR/MWh. There are also individual days where the price difference between zones is more than 10 EUR/MWh, but on average the differences are small. [^7] In 2018, Austria was split from the joint price zone with Germany. Demand and supply is significantly different between both countries, yet prices (after the gas crises in 2021 and 2022) are often remarkably close as well (see https://www.energy-charts.info/). The differences in the BZR study between the zones are higher (4-8 EUR/MWh), but these differences are still small considering average prices of 45-50 EUR/MWh, and compared to some earlier studies based on different data sets.
The BZR study finds significant 50% savings in redispatch costs in the CE region after switching from 1 to 2 zones, but no significant savings when moving from 2 to 3, 4, or 5 zones. We do not find these savings in our study when introducing 2 zones. However, similar to the BZR study, the savings in redispatch costs when moving from 2 to 5 zones are negligible if they exist. The BZR does not reveal all the details of their redispatch computation, but they allow for EU-wide redispatch, which is not current practice and one important difference to our study.
The total cost savings (including redispatch costs) of a zonal split (2-5 zones) are small for Germany in our study, and small for the CE region in the BZR study ($<$1%) . However, we find that the total cost savings of nodal pricing in Germany on the same data set are around 7-8%. This is remarkable, as we do not consider demand response, which is a likely consequence of nodal pricing. The savings in total costs are due to the fact that redispatch is largely avoided in nodal pricing systems. Redispatch costs were at 2.7 billion EUR in 2024 and they are expected to increase six-fold in the EU by 2040 (Thomassen et al., , 2024).
Our study suggests that very small price zones or nodal prices should be considered for Germany in the future, a pricing system that has not been considered in the EU so far. A split in 2-5 larger price zones seems not to have much effect based on the BZR data. Larger effects on welfare and redispatch costs only emerge with much smaller or nodal zones. Similar findings are made in a recent study by Gerhardt et al., (2025) based on a different proprietary data set. Another aspect to consider is that it is very difficult to determine zones that are robust over time (Dobos et al., , 2025). Bottlenecks arise at different locations at different times, and it is hardly possible to determine zones that are robust even with historical data, when different time frames are considered.
One might argue that the BZR data set is already several years old and the energy transition has progressed since then. The share of renewables in Germany’s electricity generation has increased from 42% in 2019 to 50% in 2023 (Bundesnetzagentur, , 2024). The grid is being expanded, but also the demand grows significantly. Recent forecasts predict, in Germany, a growth from 457 TWh electricity demand in 2022888https://www.energy-charts.info/index.html?l=en&c=DE to 726 TWh by 2030.999https://www.renewable-ei.org/pdfdownload/activities/02_DimitriPescia_231128.pdf? This makes it even harder to assess the impact of a zonal split. Yet, the BZR data set is one of the most accurate descriptions of the EU power system and it allows us to assess at least the changes that a split would have had if it was initiated a few years ago. The impact on total cost savings based on this data set is smaller than what one might have expected. The costs and benefits of a zonal split deserve a thorough analysis, and this study aims to contribute to this discussion.